OpenStudy (anonymous):

Evaluate the function f(x) = 2x + 2, when x = −1. (Points : 1) 4 −4 3 0

OpenStudy (anonymous):

f(-1)=2(-1)+2 f(-1)=-2+2 f(-1)=0

OpenStudy (anonymous):

thank you do u want best response and medal